# Stack More Layers Differently: High-Rank Training Through Low-Rank Updates

### v1

@@ 5.1 Ablation studies
The warm start shows the most significant improvement, dropping perplexity by almost 10 points. To investigate whether post-warmup training contributes to the loss, we measured the perplexity of the warmed-up network, which equals $27.03$. It outperforms all low-rank methods except for our final ReLoRA recipe but still demonstrates a significant difference from the final network. This demonstrates the importance of early training, similar to the concept of the lottery ticket hypothesis with rewinding Frankle et al. (2019).
@@ 6 Conclusion
In this paper, we investigated low-rank training techniques for large transformer language models. We first examined the limitations of a simple low-rank matrix factorization (LoRA) approach and observed that it struggles to effectively train high-performing transformer models. To address this issue, we proposed a novel method called ReLoRA, which leverages the rank of sum property to train a high-rank network through multiple low-rank updates. Similar to the lottery ticket hypothesis with rewinding, ReLoRA employs a full-rank training warm start before transitioning to ReLoRA. Additionally, ReLoRA introduces a merge-and-reinit (restart) strategy, a jagged learning rate scheduler, and partial optimizer resets, which collectively enhance the efficiency of ReLoRA and bring it closer to full-rank training, particularly in large networks. ReLoRA efficiency increases with the network size making it a viable candidate for multi-billion-scale training.
We firmly believe that the development of low-rank training methods holds great promise for improving the efficiency of training large language models and neural networks in general. Furthermore, low-rank training has the potential to provide valuable insights for the advancement of deep learning theories, aiding our understanding of neural network trainability through gradient descent and their exceptional generalization capabilities in the overparametrized regime.
7 Limitations and Future Work
Scaling beyond 350M
Due to limited computational resources, our experiments were constrained to training language models with up to 350M parameters. Nonetheless, ReLoRA already demonstrates promising results at this scale. However, we anticipate its true potential will be realized in the 1B+ parameter region. Additionally, while the 350M model outperforms the Control baseline, it does not continue the trend of narrowing the gap between ReLoRA and full-rank training. We attribute this to suboptimal hyperparameter choice, which requires further investigation.
Furthermore, in 60-350M experiments, even though ReLoRA significantly reduces the number of trainable parameters, we did not observe substantial improvements in memory and computation for the networks of this size. To evaluate the efficiency of our current implementation at a larger scale, we trained the 1.3B-parameter model for a small number of iterations to estimate memory and compute improvements of ReLoRA. At this scale, we observe 30% memory consumption reduction and 52% training throughput increase. We expect to observe even bigger improvements over the full-training baseline for larger networks since the number of trainable parameters for ReLoRA, similar to LoRA, increases at a much slower rate compared to the number of frozen parameters. ReLoRA implementation could be further improved by effectively utilizing gradient checkpointing for ReLoRA layers, custom backward functions, and converting frozen model weights to int8 or int4 quantized format Dettmers et al. (2022).
Comparison to other low-rank training methods
A number of approaches to low-rank training have been explored with other model architectures in earlier work Schotthöfer et al. (2022); Sui et al. (2023); Yang et al. (2020). Two aspects set our work apart from these earlier efforts. First, the approach we propose performs high-rank updates through low-rank training. Second, our work demonstrates competitiveness of the low-rank training methods in large-scale transformer language models with 100M+ parameters.
@@ 7 Limitations and Future Work
Thanks to Google Cloud for Research Program, Eric Lehman, and Artem Krovosheev for helping with computational resources for this paper.

### v4

@@ 4.2 Ablation studies
The warm start shows the most significant improvement, dropping perplexity by almost 10 points. To investigate whether post-warmup training contributes to the loss, we measured the perplexity of the warmed-up network, which equals $27.03$. It outperforms all low-rank methods except for our final ReLoRA recipe but still demonstrates a significant difference from the final network. This demonstrates the importance of early training, similar to the concept of the lottery ticket hypothesis with rewinding Frankle et al. (2019). In our experiments, unless specified otherwise, we performed warm start for about $1/4$ of the total training updates.
5 Related work
Scaling versus Efficiency
The relationship between overparametrization and neural network trainability and generalization has been extensively studied Zhang et al. (2017); Belkin et al. (2018); Frankle and Carbin (2019); Nakkiran et al. (2019); Singh et al. (2021), yet it remains a mystery Zhang et al. (2021).
Moreover, scaling laws Kaplan et al. (2020); Ghorbani et al. (2021); Hoffmann et al. (2022) demonstrate a simple and strong power-law dependence between network size and its performance across a variety of modalities. This finding not only supports overparametrization but also encourages the training of extraordinarily resource-intensive neural networks Brown et al. (2020); Chowdhery et al. (2022); Fedus et al. (2022). Nonetheless, the Lottery Ticket Hypothesis Frankle et al. (2019) suggests that overparametrization could, in principle, be minimized.
Parameter-efficient fine-tuning
Aghajanyan et al. (2021) found that pre-training reduces the amount of change to the network required to learn a new task through fine-tuning. I.e., larger networks or networks pre-trained on more data require smaller modifications in terms of the rank of the range to learn a new task. This explains the success of parameter-efficient fine-tuning methods Lialin et al. (2023) and has also motivated the development of low-rank fine-tuning methods such as LoRA Hu et al. (2022) and Compacter mahabadi et al. (2021).
Low-rank neural network training
Training low-rank representations has been explored in the context of CNN compression, regularization, and efficient training Idelbayev and Carreira-Perpinan (2020); Jaderberg et al. (2014); Sui et al. (2023); Schotthöfer et al. (2022); Lin et al. (2020); Yuan et al. (2021); Zhao et al. (2023). However, most of these methods are either specific to CNNs, do not scale well, or have not been evaluated on large transformers Vaswani et al. (2017) with hundreds of millions of parameters, which can benefit greatly from efficient training. While transformers have been shown to have a low-rank internal dimensionality and representations Aghajanyan et al. (2021); Wang et al. (2020), the study by Bhojanapalli et al. (2020) demonstrated that the low rank of key and query projections in multi-head attention bottlenecks the performance of transformers. Our own experiments (Section 6) also demonstrate that low-rank transformers perform significantly worse compared to the full-rank baseline and ReLoRA.
@@ 6 Conclusion
In this paper, we demonstrate that parameter-efficient fine-tuning methods can be adapted for pre-training large language models. We first examined the limitations of a low-rank matrix factorization (LoRA) approach and observed that it struggles to effectively train high-performing transformer models. To address this issue, we proposed ReLoRA, which leverages the rank of sum property to train a high-rank network through multiple low-rank updates. Similar to the lottery ticket hypothesis with rewinding, ReLoRA employs a full-rank training warm start before transitioning to ReLoRA. During training, ReLoRA periodically merges its parameters into the main parameters of the network, performs optimizer reset and learning rate re-warmup.
We demonstrated that ReLoRA consistently outperforms LoRA for training large transformer models. Our largest experiment demonstrated $9\%$ wall-clock time reduction in 8xA100 setup and much larger ($20-40\%$) speed improvements on cheaper hardware. Further, our results show similar performance to regular training making ReLoRA a promising candidate for improving the efficiency of large model training. Our further studies will focus on improving ReLoRA performance, efficiency, applying it to larger models and applying it to continued pre-training of existing large language models.
@@
This paper has been a journey and we are sincerely grateful to everyone who supported us. We would like to express our gratitude to Stability.ai, Eleuther.ai, and the Google Cloud for Research Program for providing computational resources essential for this research.
Eric Lehman and Artem Krivosheev, thank you for supporting this project from the very beginning.
Special thanks to Jason Phang, Hailey Schoelkopf, Enrico Shippole, and Stella Biderman for their technical advice and assistance with computational resources. Our experiments at billion-parameter scale wouldn’t be possible without your support.
This work was funded in part by an Amazon Alexa AI research award to Anna Rumshisky.
Appendix A A Practical guide to ReLoRA
In this section, we wanted to answer most common questions on hyperparameter selection. Especially how to select ReLoRA-specific hyperparameters to reliably get better performance than LoRA. In all of our experiments, we applied LoRA/ReLoRA to all of the linear layers in the model: kqv-projection layers, FFN layers and other projections, except for logits and embeddings.
We observed that $r\in\{64,128\}$ works well for all of the networks, up to 1B. One small, but important hyperparameter change from full-rank training to ReLoRA-training that was crucial for the performance was increased learning rate. ReLoRA (and LoRA) requires $1.5-2$ times larger learning rate than regular training/fine-tuning to achieve similar performance.
When taking about ReLoRA-specific hyperparameters, we did not observe significant dependence on optimizer pruning percentage as long as it’s larger than 90%. Larger pruning rates can lead to slightly better performance at the cost of possible loss instabilities during the reset. We tested several ReLoRA reset rates with 350M and 1.3B models and found that 2K iterations reset rate performed consistently well in both pre-training and fine-tuning experiments and always led to better performance than no resets. In general, we observed good results with reset rates 2K-5K.
Appendix B ReLoRA for fine-tuning
We apply ReLoRA to fine-tune T5-base (220M parameters) and T5-large (770M parameters) on the GLUE benchmark. We use the same type of learning rate scheduler as in ReLoRA pre-training and prune 90% of the low magnitude optimizer states during each LoRA merge-and-reinit (restart). The batch size is equal to 128 examples and the learning rate is tuned (from 1e-4 to 5e-4) on each model and dataset combination. We perform additional ReLoRA ablation studies using the T5-Large model and QNLI dataset. Specifically, we explore different ReLoRA ranks, optimizer state pruning rates, and the total number of ReLoRA resets.
Method SST-2 MNLI QNLI QQP RTE STS-B MRPC CoLA Avg
Adapters 94.2 86.4 93.1 88.9 75.1 91.1 88.9 64.4 85.3
Prompt Tuning 90.3 82.5 92.5 88.5 59.5 90.1 74.6 0.0 72.2
Ladder Side Tuning 94.1 85.6 93.3 88.8 71.9 90.7 90.4 58.1 84.1
Compacter* 93.9 86.1 92.9 90.4 76.3 91.0 91.5 64.4 85.8
KronA* 94.3 86.3 93.2 90.6 77.7 91.3 92.5 63.3 86.1
Full fine-tuning* 93.6 86.2 92.8 91.7 74.8 90.1 92.7 63.4 85.7
LoRA 93.92 86.12 91.95 90.62 78.34 89.96 90.52 60.04 85.18
ReLoRA 94.15 85.96 91.68 87.2 77.74 89.88 90.03 59.92 84.57
Full fine-tuning (T5-L) 94.7 89.1 91.6 89.9 78.9 90.6 88.9 57.0 85.0
LoRA (T5-L) 95.59 89.44 93.98 91.44 85.92 90.89 92.90 63.77 87.99
ReLoRA (T5-L) 95.7 89.06 93.68 91.04 84.72 90.53 90.57 61.72 87.47
Table 8: ReLoRA for fine-tuning does not outperform LoRA. Results with and * are T5-base results from Sung et al. (2022) and Edalati et al. (2022) respectively.
ReLoRA fine-tuning ablations
Table 9 shows the results of varying ReLoRA hyperparameters. A rank of 64 seems to provide the best performance. The results indicate that the model’s performance remains largely unaffected even when 99% of the optimizer states are reset. Our analysis of the jagged cosine learning rate scheduler’s impact on classification accuracy in the QNLI dataset suggests that two resets are adequate (reset rate 4000).
Rank Acc. Pruning Acc. Reset rate #resets Acc.
16 94.05 85% 94.51 6000 1 94.38
32 94.16 92% 94.33 4000 2 94.73
64 94.55 95% 94.31 2000 5 94.34
128 94.44 99% 94.56 1000 11 94.33
Table 9: ReLoRA fine-tuning ablations. We apply ReLoRA to fine-tune T5-large on the QNLI dataset and vary LoRA rank ($r$), optimizer state pruning percentage, and reset frequency of ReLoRA. Reset rate means the number of iterations between ReLoRA resets.
Appendix C Learning curves of models pre-trained in the study
In this section we present additional training loss plots for all of the models from Table 2. 60M: Figure 6, 130M: Figure 7, 250M: Figure 8, 350M: Figure 9, 1.3B: Figure 10.
Appendix D Ranks of 130M models
Figures 11 and 12 show spectral properties for 130M model.
Appendix E Smaller warm start period
Table 2 demonstrates that ReLoRA consistently outperforms the warmed-started LoRA baseline. To provide a more contrasting example, we performed additional pre-training experiments starting from just 2K warm-started network. Figure 13 shows a significant performance gain with ReLoRA over LoRA by 1.4 ppl points (ppl 23.64 vs 25.08). While the absolute performance of ReLoRA is lower compared to full-rank training in this context, these experiments validate our initial hypothesis that LoRA restarts positively impact performance.
Figure 6: 60M experiments training loss
Figure 7: 130M experiments training loss
Figure 8: 250M experiments training loss
Figure 9: 350M experiments training loss
Figure 10: 1.3B experiments training loss
Figure 11: Singular values spectra of the weight difference between ReLoRA and LoRA at 5,000 iterations (warm start) and 20,000 iterations. ReLoRA exhibits a closer resemblance to full-rank training than to LoRA, indicating its effectiveness in approximating full-rank behavior. 130M models.
Figure 12: The number of singular values $<0.1$ in attention and FCN matrices of the learned update. 130M models.
Figure 13: ReLoRA significantly outperforms LoRA when started from an early (2K steps) checkpoint.